A passenger train and a freight train leave cities that are 300 miles apart and travel towards each other. The passenger train is traveling 15mph faster than the freight train. Find the speed of each train if they pass each other after 4 hours.
Found 2 solutions by mananth, richwmiller:
Answer by mananth(16946)

Freight train x km/h
Passenger x+ 20 km/h
..
They are moving away from each other

So add up their speed.

combined speed = x+ 1 x + 20
( 2 x + 20 )
Time = 4 hours
Distance = 300 miles
Distance = speed * time
( 2 x 20 )* 4 = 300
8.00 x + 80 = 300
8 x = 300 + -80
8 x = 220
/ 8
x= 27.5 mph
Freight train = 27.5 mph
Passenger 27.5 + 20 = 47.5 mph
m.ananth@hotmail.ca

Answer by richwmiller(17219)

They are traveling towards each other as stated.
4*(x+15)+4x=300
x=30 mph freight train
x+15=45 mph passenger train
4*30=120 freight train travels 120 miles when they meet
4*45=180 passenger train traveled 180 miles when they meet